# Antithetic Noise in Diffusion Models

### v1

@@ B.3.2 QMC Experiments
[…]
@@
B.3.3 DPS Experiment
We evaluate the confidence interval length reduction benefits of antithetic initial noise across four common image inverse problems: super-resolution, motion deblurring, Gaussian deblurring, and inpainting. For each task, the forward measurement operator is applied to the true image, and noisy observations are generated using Gaussian noise with $\sigma=0.05$. We use the official implementation of DPS [7] with the following parameters:
@@ B.3.3 DPS Experiment
  • Motion deblurring is configured using the motion_blur operator with a kernel size of 61 and intensity of 0.1, which emulates motion-induced linear blur typical of camera shake, followed by additive noise.
@@
Table 9: Comparison of AMC vs. MC across tasks in DPS
Task Metric Group Mean CI Length Effi.
Motion Deblur L1 AMC 24.6878 1.1615 1.0342
MC 24.5166 1.1833
PSNR AMC 17.9595 0.3012 1.1222
MC 18.0234 0.3380
Super-resolution L1 AMC 28.1603 1.0115 1.4075
MC 28.1357 1.2000
PSNR AMC 16.8561 0.2592 1.2045
MC 16.8644 0.3122
Gaussian Deblur L1 AMC 22.6941 0.8879 1.3113
MC 22.6600 1.0171
PSNR AMC 18.6994 0.2651 1.4364
MC 18.7102 0.3179
Inpainting L1 AMC 21.8591 0.8286 1.5388
MC 21.9821 1.0279
PSNR AMC 19.0290 0.2575 1.7066
MC 18.9807 0.3364
@@ Appendix C More visualizations
Figure 19: Prompt: "most expensive sports car"
Figure 20: Prompt: "little bunny"

### v2

@@ B.6.2 QMC Experiments
where $t_{R-1,\,1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the $t$-distribution with R1R-1 degrees of freedom. If the estimates $\hat{\mu}_{r}^{\text{QMC}}$ are normally distributed, this confidence interval has exact coverage probability $1-\alpha$. In general, the validity of Student $t$ confidence interval is justified by CLT. An extensive numerical study by L’Ecuyer et al. (2023) demonstrates that Student $t$ intervals achieve the desired coverage empirically.
@@
B.6.3 DPS Experiment Implementation
We evaluate the confidence interval length reduction benefits of antithetic initial noise across three common image inverse problems: super-resolution, Gaussian deblurring, and inpainting. For each task, the forward measurement operator is applied to the true image, and noisy observations are generated using Gaussian noise with $\sigma=0.05$. We use the official implementation of DPS (Chung et al., 2023a) with the following parameters:
@@ B.6.3 DPS Experiment Implementation
B.6.4 DDS Dataset
Data used in the DDS experiment on uncertainty quantification (Section 4) were obtained from the NYU fastMRI Initiative database (Knoll et al., 2020; Zbontar et al., 2018). A listing of NYU fastMRI investigators, subject to updates, can be found at: fastmri.med.nyu.edu. The primary goal of fastMRI is to test whether machine learning can aid in the reconstruction of medical images.
@@ Appendix C More visualizations
Figure 21: CelebA-HQ Image Generated
Figure 22: DiT Class 974: geyser
Figure 23: DiT Class 387: lesser panda, red panda
Figure 24: DiT Clas 979: valley, vale
Figure 25: DiT Class 388: giant panda, panda
Figure 26: Prompt: “most expensive sports car”
Figure 27: Partial Negation of LSUN-Church
Figure 28: Partial Negation of LSUN-Bedroom